OpenStudy (anonymous):

Average Velocity: You drop a ball from a cliff above the sea. It's height above the ocean, in feet, t seconds after you throw it is s(t)=-16t^2+16t+320 Find the average velocity of the ball from time t=2 to time t=2+h

OpenStudy (anonymous):

\[\frac{ s(t+h)-s(t) }{(h+t)-t }=\frac{ s(t+h)-s(t) }{ h }\]

OpenStudy (anonymous):

Do I just solve for h? If that's the case I got -1ft/s

OpenStudy (anonymous):

I'm tempted to take the limit of s(t) as h->0 but that's just for instantaneous right?

OpenStudy (anonymous):

you are required to find an expression not a number

OpenStudy (anonymous):

simplify it to its lowest term and leave that as an aswer